write the ordinary differential equation (1+sin y)dx = (2ycosy-x(secy-tan y))dy


1
Expert's answer
2021-11-26T08:13:37-0500

(1+sin y)dx = (2ycosy-x(secy+tan y))dy

=> "\frac{dx}{dy} +x \frac{secy+tany}{1+siny}= \frac{2ycosy}{1+siny}"

=> "\frac{dx}{dy} +x \frac{(1+siny)}{cosy(1+siny)}= \frac{2ycosy}{1+siny}"

=> "\frac{dx}{dy} +x secy= \frac{2ycosy}{1+siny}"

This is a linear differential equation

Integrating factor (I.F.) = "e^{\int (secy )dy}= e^{ln(secy +tany)}"

= secy + tany = "\frac{1+siny}{cosy}"

Multiplying both sides by I.F. we get

"\frac{1+siny}{cosy}.\frac{dx}{dy} +x \frac{(secy+tany}{1+siny}.\frac{1+siny}{cosy}= \frac{2ycosy}{1+siny}.\frac{1+siny}{cosy}"

=> d("x\frac{1+siny}{cosy}" )= 2ydy

Integrating we get

"\int d( x\frac{1+siny}{cosy})= \int2ydy+C"

=> "x\frac{1+siny}{cosy}= y² + C" , where "C" is integration constant.

This is the solution of the differential equation given.